OpenStudy (anonymous):

18*2 over 9squared * 1 over 2 cubed

OpenStudy (anonymous):

do you mean : ((18*2)/sqrt(9))*1/(2^2) If so the result is 3

OpenStudy (unklerhaukus):

\[(18 \times 2) / 9^2 \times(1/2^3)\] =\[36/ 81 \times1/8\] =\[1/18\]

jhonyy9 (jhonyy9):

18*2 1 9*2*2*1 1 ----- * ------ = ---------- = ----- = 1/18 9*9 2*2*2 9*9*2*2*2 9*2